OpenStudy (anonymous):

PLEASE HELP A circle is inscribed in a square whose vertices have coordinates R(0, 4), S(6, 2), T(4, -4), and U(-2, -2). Find the equation of the circle.

OpenStudy (anonymous):

answer choices are (x - 2)² + y² = 68 (x + 2)² + y² = 20 (x - 2)² + y² = 10

OpenStudy (yrelhan4):

One of the easier ways is you find the distance between the vertices, that will be the diameter.And you find the middle point of the two vertices, that will be the centre. And then compare it with the given options.